OpenStudy (anonymous):

Simplify: x^2-9 (over) / x^2+6x+9

OpenStudy (anonymous):

\[\frac{ x ^{2}-9 }{x ^{2}+6x+9 }\]

OpenStudy (anonymous):

factor both polynomials

OpenStudy (anonymous):

I'm not good at factoring is my problem .. @Peter14

OpenStudy (anonymous):

what multiplies to -9 and adds to zero?

OpenStudy (anonymous):

-3 and 3? @Peter14

hartnn (hartnn):

yes, so x^2 -9 = (x+3)(x-3) got this ?

OpenStudy (anonymous):

Oh okay, yeah that makes sense. So then what about the bottom?

hartnn (hartnn):

or you can use the formula \(a^2-b^2= (a+b)(a-b)\) a= x and b= 3 here

hartnn (hartnn):

ok, for x^2+6x+9 you need to find 2 numbers whose product is 9 and sum is 6

OpenStudy (anonymous):

3 & 3?

hartnn (hartnn):

yes. so, i split 6x as 3x+3x x^2+6x+9 = x^2+3x+3x+ 9 now what can you factor out from 1st 2 terms x^2+3x .... ? now what can you factor out from last 2 terms 3x+9 ..... ?

OpenStudy (anonymous):

You can factor out the x & then the 3 ?

OpenStudy (anonymous):

X in the first, 3 in the second.

OpenStudy (anonymous):

Wait so would the bottom look like \[x^{2}(x+3)(x+3)+9\]

hartnn (hartnn):

more like \(\large x^2+3x+3x+9= x(x+3)+3(x+3)\)

OpenStudy (anonymous):

Okay

hartnn (hartnn):

then you can factor out x+3 entirely \(\large x^2+3x+3x+9= x(x+3)+3(x+3)= (x+3)(x+3)\) go this? what gets cancelled ?

OpenStudy (anonymous):

So I have this so far correct?: \[\frac{(x+3)(x-3) }{x(x+3)+3(x+3)}\]

hartnn (hartnn):

denominator = (x+3)(x+3) by factoring out x+3 entirely

hartnn (hartnn):

and then something gets cancelled!

OpenStudy (anonymous):

x+3 gets cancelled then right?

hartnn (hartnn):

yes! :)

hartnn (hartnn):

what remains is your answer...

OpenStudy (anonymous):

So then I'm left with: \[\frac{ x-3 }{ x+3 }\]?

hartnn (hartnn):

correct :) thats exactly your answer.

OpenStudy (anonymous):

Thank you so much!

hartnn (hartnn):

you are always welcome ^_^